# BALR-SAM: Boundary-Aware Low-Rank Adaptation of SAM for Resource-Efficient Medical Image Segmentation

### v1

@@ Abstract
Vision foundation models like the Segment Anything Model (SAM), pretrained on large-scale natural image datasets, often struggle in medical image segmentation due to a lack of domain-specific adaptation. In clinical practice, fine-tuning such models efficiently for medical downstream tasks with minimal resource demands, while maintaining strong performance, is challenging. To address these issues, we propose BALR-SAM, a boundary-aware low-rank adaptation framework that enhances SAM for medical imaging. It combines three tailored components: (1) a Complementary Detail Enhancement Network (CDEN) using depthwise separable convolutions and multi-scale fusion to capture boundary-sensitive features essential for accurate segmentation; (2) low-rank adapters integrated into SAM’s Vision Transformer blocks to optimize feature representation and attention for medical contexts, while simultaneously significantly reducing the parameter space; and (3) a low-rank tensor attention mechanism in the mask decoder, cutting memory usage by 75% and boosting inference speed. Experiments on standard medical segmentation datasets show that BALR-SAM, without requiring prompts, outperforms several state-of-the-art (SOTA) methods, including fully fine-tuned MedSAM, while updating just 1.8% (11.7M) of its parameters. Our code will be open-sourced upon acceptance.
@@ 1 Introduction
Medical image segmentation plays a crucial role in clinical workflows, which have attracted widespread attention from both industry and academia. However, achieving reliable segmentation in clinical environments is challenging due to several factors: ambiguous tissue boundaries, complex anatomical structures, and the limited availability of annotated data, which is costly and requires specialized expertise for labeling. Meanwhile, resource constraints in healthcare settings, such as hospitals with limited computational infrastructure, pose a distinct yet equally significant challenge [29, 20]. As a result, models that prioritize high diagnostic reliability, minimal parameters, fast inference, and low memory consumption are in high demand. To address this problem, Zhu et al. [31] argue that adapting a generalized, large-scale pretrained model with minimal fine-tuning for diverse medical tasks provides greater practical value than training multiple task-specific models. Xiang et al. [26] propose a unified perspective called MMET for more robust visual-semantic embedding learning on generalizable image retrieval, which can further boost the performance of generalizable person re-identification. Although adopting a base model with domain-specific adaptations can align well with clinical needs for both efficiency and scalability, these approaches always fail to achieve satisfactory performance in certain specific fields.
Recently, the Segment Anything Model (SAM) [12], pretrained on vast datasets of natural images, offers a promising foundation for such adaptation. Its ability to generalize across different segmentation tasks, supported by extensive pretraining, could help mitigate the scarcity of annotated medical data while leveraging its powerful feature extraction capabilities. However, SAM’s performance in the context of medical imaging remains suboptimal without domain-specific tuning. Designed primarily for natural objects, SAM struggles with the subtle boundaries and complex textures inherent in anatomical structures, often leading to imprecise delineations that undermine its clinical utility. Moreover, its high computational overhead—manifested in significant memory requirements and slower inference speeds—clashes with the resource-efficient demands of real-world healthcare environments.
@@
To address these limitations, we creatively propose a novel architecture named “BALR-SAM”, a comprehensive framework that synergistically integrates three novel components to enhance SAM’s capabilities for medical image segmentation while maintaining computational efficiency. First, we introduce a Complementary Detail Enhancement Network (CDEN) that leverages depthwise separable convolutions with dual-branch processing to capture fine-grained anatomical boundaries at multiple scales—a critical factor in medical imaging where tissue interface delineation directly impacts diagnostic accuracy. Second, we implement strategically designed and positioned low-rank decomposition adapters within SAM’s Vision Transformer blocks, reducing the parameter space by 94% compared to full-rank alternatives while enabling precise domain adaptation. Third, we develop a low-rank tensor attention mechanism that fundamentally transforms the mask decoder’s attention mechanism, reducing memory usage by approximately 75% during training compared to SAM’s original multi-head self-attention. Our extensive experiments on the standard medical segmentation dataset demonstrate that BALR-SAM achieves state-of-the-art performance while updating just 1.8% (11.7M) of its parameters.
2 Method
@@
The framework of BALR-SAM is shown in Fig 1. In this section, we first introduce the Complementary Detail Enhancement Network (CDEN) for generating boundary-sensitive features, followed by the Low-Rank Decomposition Adapters applied to the image encoder. Finally, we describe the Low-Rank Tensor Attention mechanism used in the mask decoder.
2.1 Complementary Detail Enhancement Network (CDEN)
Accurate boundary delineation in medical image segmentation is critical for identifying subtle anatomical structures, such as tumor margins or organ edges, which directly influence diagnostic precision. However, the Segment Anything Model (SAM) encoder, pretrained on natural images, often fails to capture these fine details in medical scans. To address this, we introduce the Complementary Detail Enhancement Network (CDEN), a lightweight module that enhances SAM’s encoder with boundary-sensitive features. CDEN leverages an efficient, multi-branch architecture to balance segmentation accuracy with the computational constraints typical of clinical environments. To be more specific, CDEN processes input images through five encoding stages, progressively reducing resolution to extract multi-scale features. The first stage (E1) preserves high-resolution spatial details—essential for outlining ambiguous boundaries, such as those in dermoscopic lesions or abdominal CT scans. To refine these features, we employ depthwise separable convolutions (DSC) [5], which reduce computational complexity by separating spatial and channel operations, making CDEN viable for resource-limited settings like small hospitals. Specifically, E1’s output is processed with two DSC layers to emphasize fine edges often lost in deeper stages.
@@ 2.1 Complementary Detail Enhancement Network (CDEN)
2.2 Low-Rank Decomposition Adapters for Image Encoder
@@ 3.1 Dataset and Implementation Details
3.2 Comparison with State-of-the-art Methods
Multi-organ and skin lesion segmentation In this section, we compare our proposed method with state-of-the-art medical image segmentation approaches, including both classical segmentation models and recent SAM-based automatic (prompt-free) methods. Table 1 presents the comprehensive comparison results on both the ISIC17 and Synapse datasets. Our experiments reveal several important findings. First, our parameter-efficient adaptation approach significantly outperforms the fully fine-tuned SAM, despite using only 4.2% (11.7M) of SAM’s trainable parameters. More impressively, our method surpasses even the fully fine-tuned MedSAM, which was specifically designed for medical imaging applications. This superior performance is consistent across both datasets, with particularly notable improvements in boundary precision and complex anatomical structure delineation.
Polyp segmentation To comprehensively evaluate the effectiveness of our model in polyp segmentation, we compare the proposed BALR-SAM method with several existing polyp segmentation models: UNet [19], PraNet [9], ResUNet [6], SANet [28] and UM-Net [8] on CVC-300 [24], CVC-ColonDB [23] and ETIS [22] datasets, qualitative analysis results are presented in Table 2. It can be easily observed that the proposed BALR-SAM model achieves the highest mDice across the three unseen datasets, demonstrating strong generalization performance. Notably, on the ETIS dataset, which differs significantly from the training set, our BALR-SAM model attains an Accuracy of 98.91 and Dice of 79.30, surpassing the second-ranked UM-Net [8] model by +1.96% and +14.82%, respectively.
@@ 3.2 Comparison with State-of-the-art Methods
Model Params Skin Synapse ISIC17 Aorta GB KL KR Liver PC SP SM Avg UNet [19] 34.53M 83.07 84.00 56.70 72.41 62.64 86.98 48.73 81.48 67.96 70.11 AttnUNet [15] 34.88M 83.66 82.61 61.94 76.07 70.42 87.54 46.70 80.67 67.66 71.70 PolypPVt [7] 25.11M 85.56 82.34 66.14 81.21 73.78 94.37 59.34 88.05 79.40 78.08 SwinUNet [4] 27.17M 83.97 81.76 65.95 82.32 79.22 93.73 53.81 88.04 75.79 77.58 PVT-CASCADE [17] 34.12M 85.50 83.01 70.59 82.23 80.37 94.08 64.43 90.10 83.69 81.06 nnUNet [11] 16.00M 90.80 93.01 71.77 85.57 88.18 97.23 83.01 91.86 85.26 86.99 VM-UNet [21] 27.43M 89.03 86.40 69.41 86.16 82.76 94.17 58.80 89.51 81.40 81.08 EMCAD [18] 26.76M 85.95 88.14 68.87 88.08 84.10 95.26 68.51 92.17 83.92 83.63 SAM [12] 636.00M 81.6 82.41 59.63 80.24 69.48 83.47 57.28 78.61 61.27 71.55 AI-SAM [16] - 86.6 88.89 74.53 86.56 85.01 96.30 72.84 90.32 79.24 84.21 SAMed [27] 18.81M 84.42 87.77 69.11 80.45 79.95 94.80 72.17 88.72 82.06 81.88 MedSAM [14] 636.00M 85.80 84.52 67.73 83.50 81.35 89.23 64.72 87.13 72.14 78.79 BALR-SAM (Ours) 11.7M 91.14 90.47 76.97 90.40 88.71 96.74 83.79 91.88 84.73 87.96 P-values < 1e-2 (DSC)
@@
3.3 Ablation Studies
@@
  $\times$   $\times$   $\times$   14.3   21.4
  ✓   $\times$   $\times$   64.8   74.6
  ✓   ✓   $\times$   77.2   85.6
  ✓   ✓   ✓   84.1   91.1
@@ 3.4 Visual results
To visually demonstrate the effectiveness of our proposed method, we provide a visualization of the segmentation results for polyp segmentation in Figure 2. To be more specific, the visual results of the traditional methods in Figure 2 reveals various segmentation errors, including redundancy, speckles, and omissions across different methods. On the contrary, the segmentation results of the proposed BALR-SAM method exhibit the highest overlap with the GT areas, closely resembling the ground truth. In the second and third rows of Figure 2, the polyp tissues are relatively large and exhibit minimal differences from the surrounding healthy tissues, posing significant challenges for precise segmentation methods. Our BALR-SAM method utilize Vision Transformer (ViT) as the network backbone, enabling a multi-level representation of image information, which results in better segmentation outcomes. For the small polyps in the first and fourth rows, our method employs the Complementary Detail Enhancement block to enhance the interaction between global and local information, allowing for more accurate identification and segmentation of small targets. The segmentation results in the fifth row demonstrate that our method can accurately segment the edges of polyps compared to other methods. This is attributed to the design of the SAM module in our proposed method, which selectively aggregates detail and semantic features, effectively improving the edge detection accuracy. In addition, similar improvements can also be easily observed in Figure 3 for skin lesion segmentation task.
@@ 4 Conclusion
Acknowledgement
This work was partially supported by the National Natural Science Foundation of China under Grant No.62301315, Startup Fund for Young Faculty at SJTU (SFYF at SJTU) under Grant No.23X010501967, Shanghai Municipal Health Commission Health Industry Clinical Research Special Project under Grant No.202340010 and 2025 Key Research Initiatives of Yunnan Erhai Lake National Ecosystem Field Observation Station under Grant No.2025ZD03. The authors would like to thank the anonymous reviewers for their valuable suggestions and constructive criticisms.

### v2

@@ Abstract
Vision foundation models like the Segment Anything Model (SAM), pretrained on large-scale natural image datasets, often struggle in medical image segmentation due to a lack of domain-specific adaptation. In clinical practice, fine-tuning such models efficiently for medical downstream tasks with minimal resource demands, while maintaining strong performance, is challenging. To address these issues, we propose BALR-SAM, a boundary-aware low-rank adaptation framework that enhances SAM for medical imaging. It combines three tailored components: (1) a Complementary Detail Enhancement Network (CDEN) using depth-wise separable convolutions and multi-scale fusion to capture boundary-sensitive features essential for accurate segmentation; (2) low-rank adapters integrated into SAM’s Vision Transformer blocks to optimize feature representation and attention for medical contexts, while simultaneously significantly reducing the parameter space; and (3) a low-rank tensor attention mechanism in the mask decoder, cutting memory usage by 75% and boosting inference speed. Experiments on standard medical segmentation datasets show that BALR-SAM, without requiring prompts, outperforms several state-of-the-art (SOTA) methods, including fully fine-tuned MedSAM, while updating just 1.8% (11.7M) of its parameters. The code is publicly available at https://github.com/Misterursw/cad.
keywords:
Segment Anything Model , Medical Image Segmentation , Parameter Efficient Fine-tuning
journal: Nuclear Physics B
[…]
[label1]organization=Shanghai Jiao Tong University, addressline=No 800, Dongchuan Road, Minhang District, city=Shanghai, postcode=200240, state=Shanghai, country=China
@@ 1 Introduction
Medical image segmentation plays a crucial role in clinical workflows, which have attracted widespread attention from both industry and academia. However, achieving reliable segmentation in clinical environments is challenging due to several factors: ambiguous tissue boundaries, complex anatomical structures, and the limited availability of annotated data, which is costly and requires specialized expertise for labeling. Meanwhile, resource constraints in healthcare settings, such as hospitals with limited computational infrastructure, pose a distinct yet equally significant challenge [1, 2]. As a result, models that prioritize high diagnostic reliability, minimal parameters, fast inference, and low memory consumption are in high demand. To address this problem, [3] argue that adapting a generalized, large-scale pretrained model with minimal fine-tuning for diverse medical tasks provides greater practical value than training multiple task-specific models. Xiang et al. [4] propose a unified perspective called MMET for more robust visual-semantic embedding learning on generalizable image retrieval, which can further boost the performance of generalizable person re-identification. Although adopting a base model with domain-specific adaptations can align well with clinical needs for both efficiency and scalability, these approaches always fail to achieve satisfactory performance in certain specific fields.
Recently, the Segment Anything Model (SAM) [5], pretrained on vast datasets of natural images, offers a promising foundation for such adaptation. Its ability to generalize across different segmentation tasks, supported by extensive pretraining, could help mitigate the scarcity of annotated medical data while leveraging its powerful feature extraction capabilities. However, SAM’s performance in the context of medical imaging remains suboptimal without domain-specific tuning. Designed primarily for natural objects, SAM struggles with the subtle boundaries and complex textures inherent in anatomical structures, often leading to imprecise delineations that undermine its clinical utility. Moreover, its high computational overhead—manifested in significant memory requirements and slower inference speeds—clashes with the resource-efficient demands of real-world healthcare environment.
@@
To address these limitations, we creatively propose a novel architecture named “BALR-SAM”, a comprehensive framework that synergistically integrates three novel components to enhance SAM’s capabilities for medical image segmentation while maintaining computational efficiency. First, we introduce a Complementary Detail Enhancement Network (CDEN) that leverages depth-wise separable convolutions with dual-branch processing to capture fine-grained anatomical boundaries at multiple scales—a critical factor in medical imaging. Second, we implement strategically designed and positioned low-rank decomposition adapters within SAM’s Vision Transformer blocks, reducing the parameter space by 94% compared to full-rank alternatives while enabling precise domain adaptation. Third, we develop a low-rank tensor attention mechanism that fundamentally transforms the mask decoder’s attention mechanism, reducing memory usage by approximately 75% during training compared to SAM’s original multi-head self-attention. Our extensive experiments on the standard medical segmentation dataset demonstrate that BALR-SAM could achieve state-of-the-art performance while updating just 1.8% (11.7M) of its parameters.
The remainder of this paper is structured as follows. In Section 2, we give the related works based on hand-crafted based approaches and deep learning based methods in medical area, and then briefly introduce our method. In Section 3, the details of our Complementary Detail Enhancement Network, Low-Rank Decomposition Adapters, as well as the Low-Rank Tensor Attention Mechanism, are presented. Extensive evaluations compared with state-of-the-art methods and comprehensive analyses of the proposed approach are reported in Section 4. Finally, conclusion of this paper and discussion of future works are presented in Section 5.
2 Related Work
In this section, we give a brief review of the related works on common medical image segmentation methods. The core idea of these existing methods is to efficiently model contextual information. These methods can be roughly divided into CNN-based approaches, Transformer-based approaches and SSM-based approaches.
2.1 CNN-based Approaches
Medical image segmentation methodologies have evolved through several distinct paradigms. Initially, Convolutional Neural Networks (CNNs) revolutionized the field. The Fully Convolutional Network [11] first introduced end-to-end architectures with skip connections. Subsequently, the U-Net [12] established the dominant encoder-decoder structure, which effectively captured multi-scale context while preserving localization detail, achieving strong performance even with limited data. This success spurred numerous variants. For instance, UNet++ [13] utilized nested and dense skip connections to bridge the semantic gap between the encoder and decoder. Concurrently, Attention U-Net [14] integrated attention gates to refine feature maps by suppressing irrelevant regions. Moreover, 3D U-Net [15] directly extended the architecture for volumetric data analysis. Beyond U-Net, other foundational CNN designs were adopted; ResNet [16] enabled deeper networks through residual learning, while DeepLabv3+ [17] employed atrous convolutions for multi-scale object segmentation. Additionally, attention modules like CBAM [18] were incorporated to enhance feature representation. Nevertheless, CNNs possess an inherent limitation in modeling long-range dependencies due to their localized receptive fields, which motivated the exploration of alternative architectures.
2.2 Transformer-based Approaches
Transformers emerged as a powerful alternative, adept at capturing global contextual relationships. The Vision Transformer (ViT) [19] first demonstrated applying a pure transformer to sequences of image patches. Following this, the Swin Transformer [20] introduced a hierarchical structure and an efficient shifted-window self-attention mechanism. These architectures were quickly adapted for medical imaging. TransUNet [21], for example, combined a transformer encoder with a CNN decoder to leverage both global context and precise localization. Guo et al. [22] propose UCTNet where transformers focus on establishing global dependency for CNN’s unreliable regions. Similarly, UNETR [23] utilized a pure transformer encoder for 3D volumetric segmentation. Further refinements led to models like Swin UNETR [24], which leveraged the Swin backbone, often with self-supervised pre-training, for state-of-the-art 3D segmentation. More recently, UNETR++ [25] proposed more efficient attention blocks to reduce the high computational cost associated with transformers. However, the quadratic complexity and substantial data requirements of transformers, especially for 3D volumes, remained significant challenges.
2.3 SSM-based Approaches
To address these efficiency concerns, State Space Models (SSMs) have recently gained prominence. SSMs offer a compelling approach to model long-range dependencies with linear computational complexity. The S4 model [26] provided the theoretical foundation, which was later enhanced by Mamba [27] through a selective scan mechanism that enables content-based reasoning. These principles were translated to vision tasks in models such as Vision Mamba (Vim) [28] and VMamba [29], which introduced 2D scanning mechanisms. In medical segmentation, U-Mamba [30] integrated CNNs for local feature extraction with SSMs for global context modeling. In contrast, VM-UNet [31] proposed a pure SSM-based U-shaped architecture. Likewise, Mamba-UNet [32] synergized the U-Net framework with Mamba’s efficient sequence modeling. SSMs are particularly advantageous for 3D data; for instance, SegMamba [33] was specifically designed with a tri-oriented Mamba module for 3D volumetric segmentation. Consequently, SSMs present a highly efficient alternative to transformers, though they are a nascent technology with architectural best practices still rapidly evolving.
3 Our Method
In this section, we first introduce the Complementary Detail Enhancement Network (CDEN) for generating boundary-sensitive features, followed by the Low-Rank Decomposition Adapters applied to the image encoder. Finally, we describe the Low-Rank Tensor Attention mechanism used in the mask decoder. The framework of proposed BALR-SAM is shown in Figure 1.
@@ 3 Our Method
3.1 Complementary Detail Enhancement Network
Accurate boundary delineation in medical image segmentation is critical for identifying subtle anatomical structures, such as tumor margins or organ edges, which directly influence diagnostic precision. However, the Segment Anything Model (SAM) encoder, pretrained on natural images, often fails to capture these fine details in medical scans. To address this, we introduce the Complementary Detail Enhancement Network (CDEN), a lightweight module that enhances SAM’s encoder with boundary-sensitive features. CDEN leverages an efficient, multi-branch architecture to balance segmentation accuracy with the computational constraints typical of clinical environments. To be more specific, CDEN processes input images through five encoding stages, progressively reducing resolution to extract multi-scale features. The first stage (E1) preserves high-resolution spatial details—essential for outlining ambiguous boundaries, such as those in dermoscopic lesions or abdominal CT scans. To refine these features, we employ depth-wise separable convolutions (DSC) [34], which reduce computational complexity by separating spatial and channel operations, making CDEN viable for resource-limited settings like small hospitals. Specifically, E1’s output is processed with two DSC layers to emphasize fine edges often lost in deeper stages.
@@ 3.1 Complementary Detail Enhancement Network
3.2 Low-Rank Decomposition Adapters
@@ 4.1 Dataset and Implementation Details
Methods Params ISIC17 Synapse
DSC Aorta GB KL KR Liver PC SP SM Avg
UNet [41] 34.53M 83.07 84.00 56.70 72.41 62.64 86.98 48.73 81.48 67.96 70.11
AttnUNet [42] 34.88M 83.66 82.61 61.94 76.07 70.42 87.54 46.70 80.67 67.66 71.70
PolypPVt [43] 25.11M 85.56 82.34 66.14 81.21 73.78 94.37 59.34 88.05 79.40 78.08
SwinUNet [44] 27.17M 83.97 81.76 65.95 82.32 79.22 93.73 53.81 88.04 75.79 77.58
nnUNet [45] 16.00M 90.80 93.01 71.77 85.57 88.18 97.23 83.01 91.86 85.26 86.99
VM-UNet [46] 27.43M 89.03 86.40 69.41 86.16 82.76 94.17 58.80 89.51 81.40 81.08
EMCAD [47] 26.76M 85.95 88.14 68.87 88.08 84.10 95.26 68.51 92.17 83.92 83.63
SAM [5] 636.00M 81.6 82.41 59.63 80.24 69.48 83.47 57.28 78.61 61.27 71.55
AI-SAM [10] - 86.6 88.89 74.53 86.56 85.01 96.30 72.84 90.32 79.24 84.21
SAMed [48] 18.81M 84.42 87.77 69.11 80.45 79.95 94.80 72.17 88.72 82.06 81.88
MedSAM [49] 636.00M 85.80 84.52 67.73 83.50 81.35 89.23 64.72 87.13 72.14 78.79
BALR-SAM (Ours) 11.7M 91.14 90.47 76.97 90.40 88.71 96.74 83.79 91.88 84.73 87.96
P-values << 1e-2 (DSC)
@@
4.2 Comparison with State-of-the-art Methods
Multi-Organ and Skin Lesion Segmentation. In this section, we compare our proposed method with state-of-the-art medical image segmentation approaches, including both classical segmentation models and recent SAM-based automatic (prompt-free) methods. Table 1 presents the comprehensive comparison results on both the ISIC17 and Synapse datasets. Our experiments reveal several important findings. First, our parameter-efficient adaptation approach significantly outperforms the fully fine-tuned SAM, despite using only 4.2% (11.7M) of SAM’s trainable parameters. More impressively, our method surpasses even the fully fine-tuned MedSAM, which was specifically designed for medical imaging applications. This superior performance is consistent across both datasets, with particularly notable improvements in boundary precision and complex anatomical structure delineation.
Polyp Segmentation. To comprehensively evaluate the effectiveness of our model in polyp segmentation, we compare the proposed BALR-SAM method with several existing polyp segmentation models: UNet [41], PraNet [51], ResUNet [50], SANet [52] and UM-Net [53] on CVC-300 [38], CVC-ColonDB [39] and ETIS [40] datasets, qualitative analysis results are presented in Table 2. It can be easily observed that the proposed BALR-SAM model achieves the highest mDice across the three unseen datasets, demonstrating strong generalization performance. Notably, on the ETIS dataset, which differs significantly from the training set, our BALR-SAM model attains an Accuracy of 98.91 and Dice of 79.30, surpassing the second-ranked UM-Net [53] model by +1.96% and +14.82%, respectively.
@@ 4.2 Comparison with State-of-the-art Methods
14.3 21.4
64.8 74.6
77.2 85.6
84.1 91.1
@@ 4.4 Visual results
To visually demonstrate the effectiveness of our proposed method, we provide a visualization of the segmentation results for polyp segmentation in Figure 2. To be more specific, the visual results of the traditional methods in Figure 2 reveals various segmentation errors, including redundancy, speckles, and omissions across different methods. On the contrary, the segmentation results of the proposed BALR-SAM method exhibit the highest overlap with the GT areas, closely resembling the ground truth. In the second and third rows of Figure 2, the polyp tissues are relatively large and exhibit minimal differences from the surrounding healthy tissues, posing significant challenges for precise segmentation methods. Our BALR-SAM method utilizes Vision Transformer (ViT) as the network backbone, enabling a multi-level representation of image information, which results in better segmentation outcomes. For the small polyps in the first and fourth rows, our method employs the Complementary Detail Enhancement block to enhance the interaction between global and local information, allowing for more accurate identification and segmentation of small targets. The segmentation results in the fifth row demonstrate that our method can accurately segment the edges of polyps compared to other methods. This is attributed to the design of the SAM module in our proposed method, which selectively aggregates detail and semantic features, effectively improving the edge detection accuracy. In addition, similar improvements can also be easily observed in Figure 3 for skin lesion segmentation task.